# Study of the mass of pseudoscalar glueball with a deep neural network

### v1

@@ I Introduction
Understanding glueballs is crucial for deepening our comprehension of non-perturbative QCD effects and the confinement phenomenon.The exploration of glueballs offers profound insights into the non-perturbative domain of QCD. This paper aims to study the mass of pseudoscalar glueball. To achieve this, it is necessary to introduce the topological charge density[2]
@@
In calculating the correlation function of the topological charge density, I need to use lattice QCD. Consequently, it is necessary to discretize certain physical quantities. In this paper, I choose the clover improved lattice discretization of the field strength tensor $F_{\mu\nu}\left(x\right)$ and it can be written as[2, 4]
@@ III.1 Structure
Figure 1: The structure of the DNN used to obtain the mass of the pseudoscalar glueball.
In the application phase, the inputs and outputs of the DNN are explained as follows. The input is a sequence $\left(c_{1},c_{2},\cdots,c_{N}\right)$ composed of the negative values $\widetilde{C}_{q}\left(r_{n}\right)$ of the correlation function calculated from lattice QCD at different positions. The value of $N$ varies for different lattice sizes, and in this paper, $N$ is 35. The output consists of two parameters, $P_{1}$ and $P_{2}$, both within the range (0, 1). The Sigmoid function in the last layer ensures that the output is within this range. These output parameters, $P_{1}$ and $P_{2}$, are proportional to the parameters $A$ and $m$ in $\widetilde{C}_{q}\left(r_{n}\right)$. Since the range of the DNN’s output is (0, 1), the mass $m$ of the pseudoscalar glueball needs to be scaled to fall within this range, corresponding to $P_{2}$. Because $aM$ is known to be a quantity between 0 and 2 [12], $aM/2$ becomes a quantity between 0 and 1. Thus, by scaling $M$ by $a/2$, it is adjusted to fall within the range of 0 to 1. Similarly, the parameter $A$ in $\widetilde{C}_{q}\left(r\right)$ is scaled accordingly. Therefore, the output parameter $P_{1}$ is proportional to the parameter $A$, and $P_{2}$ corresponds to $aM/2$. By obtaining the value of $aM/2$ from the DNN, the value of $M$ can be further determined. The detailed procedure on how the DNN is trained to ensure that $P_{1}$ and $P_{2}$ are proportional to the parameters $A$ and $m$ given the input $\left(c_{1},c_{2},\cdots,c_{N}\right)$ is elaborated in the training phase of the DNN.
The explanation for the parameter number in Tab. 2 is as follows. Taking Linear-1 as an example, the input to Linear-1 is a tensor with the shape $[-1,35]$, where $-1$ is a placeholder corresponding to the batch size. Linear-1 has 6400 neurons, and each neuron has a bias. Therefore, the Parameter number is calculated as $35\times 6400+6400=230400$.
@@ III.2 Training
Figure 2: DNN training process.
For the preparation of the training data, two values were randomly selected from a uniform distribution between 0 and 1 as $l_{1}$ and $l_{2}$. These values, along with $r_{n}$, were then used to compute $y_{n}=f(r_{n},l_{1},l_{2})$, resulting in an array $\left(y_{1},y_{2},\cdots,y_{N}\right)$, to which noise was added. A total of 10,000 such processed arrays constituted the training dataset. From this dataset, arrays with batche size = 100 were selected and input into the DNN, which was trained to produce outputs $o_{1}$ and $o_{2}$ matching $l_{1}$ and $l_{2}$. Essentially, $l_{1}$ and $l_{2}$ serve as the labels for the arrays $\left(y_{1},y_{2},\cdots,y_{N}\right)$, and the goal of the DNN is to output the correct labels given the input arrays.
@@ IV Result
For the LS method, I simulated a total of 1000 configurations and calculated the correlation function $C_{q}\left(r_{n}\right)$ from these configurations. Then the mass of the pseudoscalar glueball m=2370(92)MeV was fitted from 1000 ${\widetilde{C}}_{q}\left(r_{n}\right)$ data and details of the calculations are shown in Tab. 4, when 25 CPU cores were used in calculations.
For the DNN, models were trained with both noisy and noise-free training data. These models were then applied to $\widetilde{C}_{q}(r_{n})$ with different amounts of data to determine the mass of the pseudoscalar glueball. As shown in Fig. 4, the model trained with noisy data performed more accurately in obtaining the mass of the pseudoscalar glueball. For 1000 $\widetilde{C}_{q}(r_{n})$ data, the model trained with noisy yielded a mass of $m=2394(81)MeV$, which is consistent with results obtained by other researchers[12].
In terms of data details, I calculated the negative part $\widetilde{C}_{q}(r_{n})$ of the correlation function for every ten configurations, forming sequences $(c_{1},c_{2},\cdots,c_{N})$ at different distances $r_{n}$. Traditional LS method and ML method were then used to determine the mass of the pseudoscalar glueball, with errors handled using the jackknife method. The results are shown in Tab. 3.
@@
The important thing in the study is the accuracy of the physics results for LS and DNN. It can be observed from Tab.3 and Fig. 5 that the data error gradually decreases as the volume of data increases for both LS and DNN.In cases with the same amount of data, ML yields smaller errors in determining the mass of the pseudoscalar glueball compared to the LS. Furthermore, the fitting results obtained by ML are more stable across different data volumes. In cases with small data volume, the mass obtained by the LS is inaccurate. In contrast, the DNN can still produce good result even with small data volume.
@@
One possible explanation for these results is as follows. Fully connected layers in this DNN, by connecting all input units, can learn the global features of the input data. Despite the presence of noise in the data, the fully connected layers can still identify useful patterns. During training, the network adjusts its weights to fit the effective patterns in the data while ignoring the noisy features as much as possible. The stacking of multiple fully connected layers enhances the network’s nonlinear representation capability, allowing it to differentiate useful features from noise to some extent. Through this multi-layer structure, the network can extract higher-level features layer by layer, which helps to filter out the impact of noise at higher levels. Therefore, DNNs can suppress the influence of noise and extract effective features from the data to a certain degree. Related research also shows that DNNs can filter out noise in the data[15]. In contrast, the LS method determines model parameters by minimizing the sum of the squared errors between the original data and the predicted data, which makes it more susceptible to the impact of noise when the data volume is small, leading to instability in the results.

### v2

@@ I Introduction
Understanding glueballs is crucial for deepening our comprehension of non-perturbative QCD effects and the confinement phenomenon.The exploration of glueballs offers profound insights into the non-perturbative domain of QCD. This paper aims to study the mass of pseudoscalar glueball. To achieve this, it is necessary to introduce the topological charge density[2, 3]
@@
In calculating the correlation function of the topological charge density, I need to use lattice QCD. Consequently, it is necessary to discretize certain physical quantities. In this paper, I choose the clover improved lattice discretization of the field strength tensor $F_{\mu\nu}\left(x\right)$ and it can be written as[2]
@@ III.1 Structure
Figure 1: The structure of the DNN used to obtain the mass $m$ of the pseudoscalar glueball. In the application phase, the input is a sequence $\left(c_{1},c_{2},\cdots,c_{N}\right)$ composed of the negative values $\widetilde{C}_{q}\left(r_{n}\right)$ of the correlation function at different positions $r_{n}$ and the outputs $P_{1}$ and $P_{2}$ are proportional to the parameter $A$ and the mass $m$.
@@
The explanation for the parameter number in Tab. 2 is as follows. Taking layer Linear-1 as an example, the input to Linear-1 is a tensor with the shape $[-1,35]$, where $-1$ is a placeholder. Linear-1 has 6400 neurons, and each neuron has a bias. Therefore, the Parameter number is calculated as $35\times 6400+6400=230400$.
In the application phase, the inputs and outputs of the DNN are shown in Fig. 1. The input is a sequence $\left(c_{1},c_{2},\cdots,c_{N}\right)$ composed of the negative values $\widetilde{C}_{q}\left(r_{n}\right)$ of the correlation function calculated from lattice QCD at different positions. The value of $N$ varies for different lattice sizes, and in this paper, $N$ is 35. The output consists of two parameters, $P_{1}$ and $P_{2}$, both within the range (0, 1). The Sigmoid function in the last layer ensures that the output is within this range. These output parameters, $P_{1}$ and $P_{2}$, are proportional to the parameters $A$ and $m$ in $\widetilde{C}_{q}\left(r_{n}\right)$. Since the range of the DNN’s output is (0, 1), the mass $m$ of the pseudoscalar glueball needs to be scaled to fall within this range, corresponding to $P_{2}$. Because $aM$ is known to be a quantity between 0 and 2 [12], $aM/2$ becomes a quantity between 0 and 1. Thus, by scaling $M$ by $a/2$, it is adjusted to fall within the range of 0 to 1. Similarly, the parameter $A$ in $\widetilde{C}_{q}\left(r\right)$ is scaled accordingly. Therefore, the output parameter $P_{1}$ is proportional to the parameter $A$, and $P_{2}$ corresponds to $aM/2$. By obtaining the value of $aM/2$ from the DNN, the value of $M$ can be further determined. The detailed procedure on how the DNN is trained to ensure that $P_{1}$ and $P_{2}$ are proportional to the parameters $A$ and $m$ given the input $\left(c_{1},c_{2},\cdots,c_{N}\right)$ is elaborated in the training phase of the DNN.
@@ III.2 Training
Figure 2: DNN training process.This model is trained to recognize the parameters $l_{1}$ and $l_{2}$ of function $f(r,l_{1},l_{2})$. In other words, this neural network is equivalent to a functional $F$, which maps the input function $f(r,l_{1},l_{2})$ to its parameters $l_{1}$ and $l_{2}$ approximately. In the actual operation, the input of DNN is a sequence of $f$ with noise, and $o_{1}$ and $o_{2}$ are output corresponding to the parameters $l_{1}$ and $l_{2}$. Since $\widetilde{C}_{q}(r)$ is a special case of $f$, this functional $F$ should also be able to determine the two parameters of $\widetilde{C}_{q}(r)$, specifically $-A/100$ and $am/2$, thereby obtaining the pseudoscalar glueball mass $m$.
For the preparation of the training data, two values were randomly selected from a uniform distribution between 0 and 1 as $l_{1}$ and $l_{2}$. These values, along with $r_{n}$, were then used to compute $y_{n}=f(r_{n},l_{1},l_{2})$, resulting in an array $\left(y_{1},y_{2},\cdots,y_{N}\right)$, to which noise was added. A total of 10,000 such processed arrays constituted the training dataset. From this dataset, arrays with batch size = 100 were selected and input into the DNN, which was trained to produce outputs $o_{1}$ and $o_{2}$ matching $l_{1}$ and $l_{2}$. Essentially, $l_{1}$ and $l_{2}$ serve as the labels for the arrays $\left(y_{1},y_{2},\cdots,y_{N}\right)$, and the goal of the DNN is to output the correct labels given the input arrays.
@@
It is worth noting that the input data in the training phase and the application phase are different. The input data in the training phase is $y=\left(y_{1},y_{2},\ldots,y_{N}\right)$ or $y+noise$, and the input data in the application phase is $c=(c_{1},c_{2},\ldots,c_{N})$.
@@ IV Result
In terms of data details, I simulated a total of 1000 configurations and calculated the correlation function $C_{q}\left(r_{n}\right)$ from these configurations, when 25 CPU cores were used in calculations. I calculated the negative part $\widetilde{C}_{q}(r_{n})$ of the correlation function for every ten configurations, forming sequences $(c_{1},c_{2},\cdots,c_{N})$ at different distances $r_{n}$. Traditional LS method and ML method were then used to determine the mass of the pseudoscalar glueball, with errors handled using the jackknife method. The results are shown in Tab. 3. Both the LS method and the ML method use $\chi^{2}$ test to assess goodness-of-fit.The calculation shows that the values of $\chi^{2}$ per degree of freedom are all less than 2.5, indicating that two methods generally suit the data well.
For the LS method, the mass of the pseudoscalar glueball m=2370(92)MeV was fitted from 1000 data and details of the calculations are shown in Tab. 3.
For the DNN, models were trained with both noisy and noise-free training data. These models were then applied to $\widetilde{C}_{q}(r_{n})$ with different amounts of data to determine the mass of the pseudoscalar glueball. As shown in Fig. 4, the model trained with noisy data performed more accurately in obtaining the mass of the pseudoscalar glueball. For 1000 data, the model trained with noisy yielded a mass of $m=2394(81)MeV$, which is consistent with results obtained by other researchers[12, 15, 16].
@@
The important thing in the study is the accuracy of the physics results for LS and DNN. It can be observed from Tab.3 and Fig. 5 that the data error gradually decreases as the volume of data increases for both LS and DNN.In cases with the same amount of data, ML yields smaller errors in determining the mass of the pseudoscalar glueball compared to the LS. Furthermore, the fitting results obtained by ML are more stable across different data volumes.
@@
One possible explanation for these results is as follows. Fully connected layers in this DNN, by connecting all input units, can learn the global features of the input data. Despite the presence of noise in the data, the fully connected layers can still identify useful patterns. During training, the network adjusts its weights to fit the effective patterns in the data while ignoring the noisy features as much as possible. The stacking of multiple fully connected layers enhances the network’s nonlinear representation capability, allowing it to differentiate useful features from noise to some extent. Through this multi-layer structure, the network can extract higher-level features layer by layer, which helps to filter out the impact of noise at higher levels. Therefore, DNNs can suppress the influence of noise and extract effective features from the data to a certain degree. Related research also shows that DNNs can filter out noise in the data[17]. In contrast, the LS method determines the fitted values by minimizing the sum of the squared errors between the original data and the predicted data, which makes it more susceptible to the impact of noise when the data volume is small, leading to instability in the results.